OpenStudy (anonymous):

Find the ordered pairs for the x- and y-intercepts of the equation 3x − 2y = 18 and select the appropriate option below. The x-intercept is (0, 6); the y-intercept is (–9, 0). The x-intercept is (6, 0); the y-intercept is (0, –9). The x-intercept is (3, 0); the y-intercept is (0, –2). The x-intercept is (0, 3); the y-intercept is (–2, 0).

OpenStudy (anonymous):

ill give fan & medal

OpenStudy (anonymous):

x-intercept are in the form of : (x,0) y-intercept are in the form of: (0,y)

OpenStudy (anonymous):

3x − 2y = 18 (x,0) 3x-2(0)=18 3x=18 x=6 x-intercept= (6,0)

OpenStudy (anonymous):

can you find the y intercept now? 3x − 2y = 18 (0,y)

OpenStudy (anonymous):

yes i found it thanks

OpenStudy (anonymous):

welcome